5. Use your financial calculator to compute the monthly payments for a vehicle that costs $11,000 if you financed the entire purchase over four years at an annual interest rate of 7.75

percent. Also calculate the loan payments assuming rates of 6.75 percent and 8.75 percent. Compare the total amount spent on the vehicle under each assumption.

ACTUAL QUESTION: The total amount spent on the vehicle if financed for four years at an annual rate of 7.75 percent is $______ round to the nearest cent
Answer #1

Loan Amount = $11,000

Time Period = 4 years

Calculating Monthly Payment at 7.75%

Using TVM Calculation,

PMT = [PV = 11,000, FV = 0, N = 48, I = 0.0775/12]

PMT = $267.25

Calculating Monthly Payment at 6.75%

Using TVM Calculation,

PMT = [PV = 11,000, FV = 0, N = 48, I = 0.0675/12]

PMT = $262.13

Calculating Monthly Payment at 8.75%

Using TVM Calculation,

PMT = [PV = 11,000, FV = 0, N = 48, I = 0.0875/12]

PMT = $272.43